QuestionCoveBot:

The largest doll is 12 inches tall. The height of each doll is 7/10(fraction) inch less than the next larger doll. Write an expression involving a power for the height of the small doll. What is the height of the smallest doll?

justus:

1st doll: 12 inches 2nd doll: \[12\times \frac{ 7 }{ 10 }\] 3rd doll:\[12 \times \frac{ 7 }{ 10 }\times \frac{ 7 }{ 10}=12(\frac{ 7 }{ 10 })^2\] 4th doll: \[12(\frac{ 7 }{ 10 })^2\]\[\times \frac{ 7 }{ 10}=12(\frac{ 7 }{ 10 })^3=4.116\]

QuestionCoveBot:

What do I add to the equation?

QuestionCoveBot:

Do I just add 7/10?

justus:

The expression for the height of the smallest doll is \[12(\frac{ 7 }{ 10})^3 \]

justus:

and the height is 4.116 inches

QuestionCoveBot:

Oh, okay. Thank you.

QuestionCoveBot:

Forgot there were 4 dolls.

justus:

Lol.

justus:

I know the expression of the smallest doll because 12 is multiplied by 0.7^(n-1)

QuestionCoveBot:

Okay.

justus:

So the height of the doll can be done by 12 * 0.7^(n-1)

QuestionCoveBot:

The height of the smallest doll?

justus:

4th doll is 12*0.7^(4-1) which equals 12*0.7^(3)

justus:

12*0.7^(3) is 4.116

QuestionCoveBot:

Thank you justus.

justus:

No. As I said above the height of the smallest doll is 4.116 and I also wrote the expression along with it too which is \[12(\frac{ 7 }{ 10 })^3\]

justus:

Of course!